OpenStudy (anonymous):

cube root of 2x^8/ cube root of 16x^2

OpenStudy (anonymous):

\[\frac{\sqrt[3]{2x^{8}}}{\sqrt[3]{16x^{2}}} = \frac{\sqrt[3]{2}x^{8/3}}{\sqrt[3]{16}x^{2/3}} = \frac{x^{6/3}}{\sqrt[3]{8}} = \frac{x^2}{2} \]

OpenStudy (anonymous):

and this one \[\sqrt(32x^10)/\sqrt(2x^4)\]

OpenStudy (anonymous):

the 32x^10